OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of negative 2 times x minus 12 all divided by negative 2

mathslover (mathslover):

\[2(x-12)/(-2) = (2x-24)/(-2) = -(2x-24)/2 = (24-2x)/2 = 2(12-x)/2 = 12-x\]